The speed v in a gas can reasonably be dependent on the pressure p, the density p, and the volume V of the gas. Using dimensional analysis, determine the exponents l,m and n in the formula v=Cp^lp^mV^n


1
Expert's answer
2022-01-10T09:09:04-0500

We have:


"v=\rho^lp^mV^n,\\\space\\
[\text{m/s}]=[\text{kg/m}^3]^l·[\text{kg/(s}·\text m)]^m·[\text{m}^3]^n,\\
[\text{m}]^1[\text s]^{-1}=[\text{kg}]^{l+m}[\text{m}]^{-3l-m+3n}[\text s]^{-m}."

Comparing the powers, we see that


"l+m=0,\\
-3l-m+3n=1,\\
-m=-1.\\\space\\
l=-1,\\
m=1,\\
n=-1/3.\\\space\\
v=\dfrac{p}{\rho\sqrt[3]{V}}."